Working together Deng and Ddolly can sew a set of uniform in 2 hours. If Dolly can do it alone in 5 hours, howw long will it take Deng to do it by herslf?
Answer by ptaylor(2198)

Let x=amount of time it takes Deng working alone
Then Deng works at the rate of 1/x uniform/hr
Deng and Dolly together work at the rate of 1/2 uniform/hour
Dolly, working alone, works at the rate of 1/5 uniform/hr
So, our equation to solve is:
1/x + 1/5= 1/2 multiply each term by 10x
10+2x=5x subtract 2x from each side
10+2x-2x=5x-2x
3x=10
x=3 1/3 hours=10/3 hours----Amount of time needed for Deng to do the job working alone
CK
1/5 +(1/10/3)=1/2
1/5 +3/10=1/2
2/10+3/10=1/2
1/2=1/2
Hope this helps---ptaylor